OpenStudy (anonymous):

The angle of elevation of a tower from a distance 100m from its foot is 60 degrees. Find the height of the tower.

OpenStudy (diyadiya):

\[\tan \theta= \frac {opp.side}{adj.Side}\] \[\tan 60= \frac{h}{100}\] \[\sqrt{3}= \frac {h}{100}\] \[h=100\sqrt{3}\]

OpenStudy (anonymous):

\[100 \text{ Tan }[60{}^{\circ}]=100 \sqrt{3}=173.2 m \]

OpenStudy (anonymous):

Even I got this answer, but in book, it is given as " 100/ root 3".

OpenStudy (anonymous):

Notify the book publishers.